OpenStudy (anonymous):

Give an example of a linear equation in standard form. Change that equation into slope and intercept form. What is the slope? From left to right, is the line going up or down? What is the y-intercept? Is the y-intercept above or below the x-axis?

OpenStudy (texaschic101):

2x + 4y = 12 (standard form) 2x + 4y = 12 4y = -2x + 12 y = -2/4x + 12/4 y = -1/2x + 3 (slope-intercept form) In y = mx + b, the slope is in the m position...so the slope is -1/2. your line is going down because of your negative slope. y = -1/2x + 3....in y = mx + b, the y intercept is in the b position...so your y intercept is (0,3) the y intercept (0,3) is above the x axis.